OpenStudy (mathhelp346):

The amount of money that must be invested at the annual rate of 12% to earn $18,000 per year is: a.$1,500,000 b.$150,000 c.$15,000 d.$1,500

OpenStudy (helder_edwin):

\[ \large P\cdot12\%=18000 \]

OpenStudy (mathhelp346):

..i dont get it

OpenStudy (helder_edwin):

for every dollar u earn 12% (12cents) so if you deposit P dollar your earn \[ \large P\cdot12\%=P\cdot\frac{12}{100} \]

OpenStudy (mathhelp346):

but then what do i do

OpenStudy (helder_edwin):

solve for P \[ \large P\cdot12\%=18000 \]

OpenStudy (mathhelp346):

so i multiply both sides by 12/100?

OpenStudy (helder_edwin):

solve for P \[ \large P\cdot\frac{12}{100}=18000 \]

OpenStudy (mathhelp346):

ok

OpenStudy (helder_edwin):

multiply by 100/12

OpenStudy (mathhelp346):

thanks

OpenStudy (helder_edwin):

u r welcome

OpenStudy (mathhelp346):

so...the answer is b.$150,000 right

OpenStudy (helder_edwin):

yes

OpenStudy (mathhelp346):

ok thanks alot

OpenStudy (helder_edwin):

u r welcome